# A uniquely solvable and positivity-preserving finite difference scheme for the Flory-Huggins-Cahn-Hilliard equation with dynamical boundary condition

### v1

@@ 1 Introduction
where $F$ is a double-well energy. Assuming homogeneous Neumann boundary conditions, as is standard, the free energy of the system, namely,
@@
where $\nu$ stands for the surface chemical potential; $\Delta_{\Gamma}$ is the Laplace-Beltrami operator on the Lipschitz continuous boundary $\Gamma:=\partial\Omega$; the constant $\kappa$ is the surface diffuse interface thickness parameter; and $\varepsilon$ is an bulk diffuse interface thickness parameter. Effectively, periodic boundary conditions are employed for the surface diffusion problem. Owing to the no mass flux $\partial_{n}\mu=0$ for the bulk problem and the periodic boundary conditions for the surface problem, one will observe that no mass is exchanged between the surface and bulk regions, which leads to respective mass conservation of $\phi$ in $\Omega$ and $\psi$ on $\Gamma$. Effectively, in the Liu-Wu model, the bulk boundary condition is given by a lower dimensional CH type equation that is coupled with the surface normal derivative term $\partial_{n}\phi$.
@@
Some numerical efforts have been reported for the Cahn-Hilliard equation coupled with dynamical boundary condition, typically with polynomial energy potential. For example, a finite element numerical scheme was proposed in [29]; an energy stability is proved, and a convergence to the weak solution is established. A reaction rate dependent dynamic boundary condition was considered in [22]. Based on the stabilized linearly implicit approach, a first-order-in-time, linear and energy stable scheme was proposed in [5], and the error analysis was provided as well. In addition, a scalar auxiliary variable (SAV) approach was applied to the Cahn-Hilliard-Hele-Shaw system in [36], with the energy stability analysis theoretically justified.
@@ 3 Unique solvability and positivity preserving properties
$\quad+\varepsilon^{-1}\left(\left\langle I(\phi),1\right\rangle_{\Omega}+\frac{1}{h}\left\langle I(\phi_{B}),1\right\rangle_{\Gamma}+\frac{1}{h}\left\langle I(\phi_{T}),1\right\rangle_{\Gamma}\right)$
$\quad+\frac{\varepsilon}{2}\left\langle\phi,L_{h}\phi\right\rangle_{\Omega}+\frac{\kappa}{2h}\left\langle\phi_{B},-\Delta_{h}^{x}\phi_{B}\right\rangle_{\Gamma}+\frac{\kappa}{2h}\left\langle\phi_{T},-\Delta_{h}^{x}\phi_{T}\right\rangle_{\Gamma}$
$\quad-\varepsilon^{-1}\theta_{0}\left(\left\langle\phi^{n},\phi\right\rangle_{\Omega}+\frac{1}{h}\left\langle\phi_{B}^{n},\phi_{B}\right\rangle_{\Gamma}+\frac{1}{h}\left\langle\phi_{T}^{n},\phi_{T}\right\rangle_{\Gamma}\right),$
@@
$=\frac{1}{s}\left\langle\psi,L_{h}^{-1}(\phi-\phi^{n})\right\rangle_{\Omega}+\frac{1}{hs}\left\langle\psi_{B},(-\Delta_{h}^{x})^{-1}(\phi_{B}-\phi^{n}_{B})\right\rangle_{\Gamma}+\frac{1}{hs}\left\langle\psi_{T},(-\Delta^{x}_{h})^{-1}(\phi_{T}-\phi^{n}_{T})\right\rangle_{\Gamma}$
$\quad+\varepsilon^{-1}\left(\left\langle I^{\prime}(\phi),\psi\right\rangle_{\Omega}+\frac{1}{h}\left\langle I^{\prime}(\phi_{B}),\psi_{B}\right\rangle_{\Gamma}+\frac{1}{h}\left\langle I^{\prime}(\phi_{T}),\psi_{T}\right\rangle_{\Gamma}\right)$
@@
$\quad-\varepsilon^{-1}\theta_{0}\left(\left\langle\phi^{n},\psi\right\rangle_{\Omega}+\frac{1}{h}\left\langle\phi_{B}^{n},\psi_{B}\right\rangle_{\Gamma}+\frac{1}{h}\left\langle\phi_{T}^{n},\psi_{T}\right\rangle_{\Gamma}\right).$ (3.17)
@@
$\quad+\frac{1}{hs}(-\Delta^{x}_{h})^{-1}(\phi_{B}-\phi^{n}_{B})_{i}+\frac{1}{h}\varepsilon^{-1}\left(I^{\prime}(\phi_{B,i})-\theta_{0}\phi^{n}_{B,i}\right)-\frac{1}{h}\kappa\Delta_{h}^{x}\phi_{B,i}+\tilde{C}_{B,n},$ (3.20)
@@
$\left.\frac{\mathrm{d}^{2}F_{h}^{n}}{\mathrm{d}\tau^{2}}(\phi+\tau\psi)\right|_{\tau=0}$ $=\frac{1}{s}\left\langle\psi,L_{h}^{-1}\psi\right\rangle_{\Omega}+\frac{1}{hs}\left\langle\psi_{B},(-\Delta_{h}^{x})^{-1}\psi_{B}\right\rangle_{\Gamma}+\frac{1}{hs}\left\langle\psi_{T},(-\Delta^{x}_{h})^{-1}\psi_{T}\right\rangle_{\Gamma}$
$\quad+\varepsilon^{-1}\left(\left\langle\psi I^{\prime\prime}(\phi),\psi\right\rangle_{\Omega}+\frac{1}{h}\left\langle\psi_{B}I^{\prime\prime}(\phi_{B}),\psi_{B}\right\rangle_{\Gamma}+\frac{1}{h}\left\langle\psi_{T}I^{\prime\prime}(\phi_{T}),\psi_{T}\right\rangle_{\Gamma}\right)$
@@
$=\frac{1}{s}\left\langle\psi,L_{h}^{-1}(\phi-\phi^{n})\right\rangle_{\Omega}+\frac{1}{hs}\left\langle\psi_{B},(-\Delta_{h}^{x})^{-1}(\phi_{B}-\phi^{n}_{B})\right\rangle_{\Gamma}+\frac{1}{hs}\left\langle\psi_{T},(-\Delta^{x}_{h})^{-1}(\phi_{T}-\phi^{n}_{T})\right\rangle_{\Gamma}$
$\quad+\varepsilon^{-1}\left(\left\langle I^{\prime}(\phi),\psi\right\rangle_{\Omega}+\frac{1}{h}\left\langle I^{\prime}(\phi_{B}),\psi_{B}\right\rangle_{\Gamma}+\frac{1}{h}\left\langle I^{\prime}(\phi_{T}),\psi_{T}\right\rangle_{\Gamma}\right)$
@@
$\quad-\varepsilon^{-1}\theta_{0}\left(\left\langle\phi^{n},\psi\right\rangle_{\Omega}+\frac{1}{h}\left\langle\phi_{B}^{n},\psi_{B}\right\rangle_{\Gamma}+\frac{1}{h}\left\langle\phi_{T}^{n},\psi_{T}\right\rangle_{\Gamma}\right).$ (3.25)
@@
$\quad+\frac{1}{hs}\Big{(}(-\Delta_{h}^{x})^{-1}(\phi^{\star}_{B}-\phi^{n}_{B})_{i_{0}}-(-\Delta_{h}^{x})^{-1}(\phi^{\star}_{B}-\phi^{n}_{B})_{i_{1}}\Big{)}$
$\quad+\varepsilon^{-1}(\frac{1}{2}+h^{-1})\Big{(}\ln(1+\phi^{\star}_{i_{0},0})-\ln(1-\phi^{\star}_{i_{0},0})\Big{)}$
$\quad+\varepsilon^{-1}(\frac{1}{2}+h^{-1})\Big{(}-\ln(1+\phi^{\star}_{i_{1},0})+\ln(1-\phi^{\star}_{i_{1},0})\Big{)}$
@@
$\quad-\varepsilon^{-1}\theta_{0}(\frac{1}{2}+h^{-1})(\phi^{n}_{i_{0},0}-\phi^{n}_{i_{1},0}).$ (3.36)
In comparison with (3.27), in which the minimum value point $(i_{0},j_{0})$ is located within the interior of the grid, a distinguishing feature of the directional derivative (3.36) is the coefficient in front of the singular logarithmic terms, which is $\varepsilon^{-1}(\frac{1}{2}+h^{-1})$, instead of $\varepsilon^{-1}$ in (3.27). In either case, the corresponding coefficients are positive, and this crucial fact will play an essential role in the positivity-preserving analysis.
@@
$\frac{1/2+1/h}{\varepsilon}\left(\ln(1+\phi^{\star}_{i_{0},0})-\ln(1+\phi^{\star}_{i_{0},0})-\ln(1-\phi^{\star}_{i_{0},0})+\ln(1-\phi^{\star}_{i_{1},0})\right)$ (3.37)
$\leq\frac{1/2+1/h}{\varepsilon}\left(\ln\left(\frac{\delta}{2-\delta}\right)+\ln\left(\frac{1-\beta_{B,0}}{1+\beta_{B,0}}\right)\right)$ (3.38)
@@
$-\frac{\theta_{0}(1+2/h)}{\varepsilon}\leq\frac{\theta_{0}(1/2+1/h)}{\varepsilon}\left(\phi^{n}_{i_{0},0}-\phi^{n}_{i_{1},0}\right)\leq\frac{\theta_{0}(1+2/h)}{\varepsilon}.$ (3.43)
@@
$\frac{1}{h^{2}}\mathrm{d}_{\tau}\!\left.F_{h}^{n}(\phi^{\star}+\tau\psi)\right|_{\tau=0}\leq\frac{1/2+1/h}{\varepsilon}\ln\left(\frac{\delta}{2+\delta}\right)+C_{4},$ (3.44)
@@
$C_{4}:=\frac{1/2+1/h}{\varepsilon}\ln\left(\frac{1-\beta_{B,0}}{1+\beta_{B,0}}\right)+\frac{C_{1}+2C_{2}/h}{s}+\frac{2\varepsilon}{h^{2}}+\theta_{0}\frac{1+2/h}{\varepsilon}.$
@@
$\frac{1}{h^{2}}\mathrm{d}_{\tau}\!\left.F_{h}^{n}(\phi^{\star}+\tau\psi)\right|_{\tau=0}\leq\frac{1/2+1/h}{\varepsilon}\ln\left(\frac{\delta}{2+\delta}\right)+C_{4}<0.$ (3.45)
@@
$\quad-\frac{1}{hs}(-\Delta_{h}^{x})^{-1}(\phi_{B}^{\star}-\phi_{B}^{n})_{i_{1}}+\varepsilon^{-1}\Big{(}\ln(1+\phi^{\star}_{i_{0},j_{0}})-\ln(1-\phi^{\star}_{i_{0},j_{0}})\Big{)}$
$\quad+\varepsilon^{-1}(\frac{1}{2}+h^{-1})\Big{(}-\ln(1+\phi^{\star}_{i_{1},0})+\ln(1-\phi^{\star}_{i_{1},0})\Big{)}$
@@
$\quad-\varepsilon^{-1}\theta_{0}\phi^{n}_{i_{0},j_{0}}+\varepsilon^{-1}\theta_{0}(\frac{1}{2}+h^{-1})\phi^{n}_{i_{1},0}.$ (3.46)
@@ 5 Numerical experiments
In this section, some numerical examples will be presented to validate the proposed scheme. A combination of multigrid method and Newton iteration is applied to deal with the nonlinear equations coupled between the interior region and the boundary section. A nonlinear full approximation scheme (FAS) multigrid method is applied for solving the semi-implicit numerical scheme (2.15) – (2.19). We present the convergence test and perform some 2-D sample computation results, while a 3-D extension could be similar. The computational domain is taken as $\Omega=(0,1)^{2}$, other parameters will be specified when mentioned.

### v2

@@ 1 Introduction
where $\phi$ is the required phase separation and $\mu$ is the chemical potential. $F$ is a double-well energy and $\varepsilon$ is an interface thickness parameter. Assuming homogeneous Neumann boundary conditions, as is standard, the free energy of the system, namely,
@@
where $\Omega$ is a bounded domain with Lipschitz continuous boundary $\Gamma:=\partial\Omega$; $\nu$ stands for the surface chemical potential; $\Delta_{\Gamma}$ is the Laplace-Beltrami operator on $\Gamma$; the constant $\kappa$ is the surface diffuse interface thickness parameter; and $\varepsilon$ is an bulk diffuse interface thickness parameter. Effectively, periodic boundary conditions are employed for the surface diffusion problem. Owing to the no mass flux $\partial_{n}\mu=0$ for the bulk problem and the periodic boundary conditions for the surface problem, one will observe that no mass is exchanged between the surface and bulk regions, which leads to respective mass conservation of $\phi$ in $\Omega$ and $\psi$ on $\Gamma$. Effectively, in the Liu-Wu model, the bulk boundary condition is given by a lower dimensional CH type equation that is coupled with the surface normal derivative term $\partial_{n}\phi$.
@@
Some numerical efforts have been reported for the Cahn-Hilliard equation coupled with dynamical boundary condition, typically with polynomial energy potential. For example, a finite element numerical scheme was proposed in [32]; an energy stability is proved, and a convergence to the weak solution is established. A reaction rate dependent dynamic boundary condition was considered in [24]. Based on the stabilized linearly implicit approach, a first-order-in-time, linear and energy stable scheme was proposed in [5], and the semi-discrete error analysis was provided as well. A second order stabilized semi-implicit scheme was analyzed in [31]. In addition, a scalar auxiliary variable (SAV) approach was applied to the Cahn-Hilliard-Hele-Shaw system in [39], with the energy stability analysis theoretically justified.
@@ 3 Unique solvability and positivity preserving properties
$\quad+\varepsilon^{-1}\left(\left\langle I(\phi),1\right\rangle_{\Omega}+\frac{2}{h}\left\langle I(\phi_{B}),1\right\rangle_{\Gamma}+\frac{2}{h}\left\langle I(\phi_{T}),1\right\rangle_{\Gamma}\right)$
$\quad+\frac{\varepsilon}{2}\left\langle\phi,L_{h}\phi\right\rangle_{\Omega}+\frac{\kappa}{h}\left\langle\phi_{B},-\Delta_{h}^{x}\phi_{B}\right\rangle_{\Gamma}+\frac{\kappa}{h}\left\langle\phi_{T},-\Delta_{h}^{x}\phi_{T}\right\rangle_{\Gamma}$
$\quad-\varepsilon^{-1}\theta_{0}\left(\left\langle\phi^{n},\phi\right\rangle_{\Omega}+\frac{2}{h}\left\langle\phi_{B}^{n},\phi_{B}\right\rangle_{\Gamma}+\frac{2}{h}\left\langle\phi_{T}^{n},\phi_{T}\right\rangle_{\Gamma}\right),$
@@
$=\frac{1}{s}\left\langle\psi,L_{h}^{-1}(\phi-\phi^{n})\right\rangle_{\Omega}+\frac{2}{hs}\left\langle\psi_{B},(-\Delta_{h}^{x})^{-1}(\phi_{B}-\phi^{n}_{B})\right\rangle_{\Gamma}+\frac{2}{hs}\left\langle\psi_{T},(-\Delta^{x}_{h})^{-1}(\phi_{T}-\phi^{n}_{T})\right\rangle_{\Gamma}$
$\quad+\varepsilon^{-1}\left(\left\langle I^{\prime}(\phi),\psi\right\rangle_{\Omega}+\frac{2}{h}\left\langle I^{\prime}(\phi_{B}),\psi_{B}\right\rangle_{\Gamma}+\frac{2}{h}\left\langle I^{\prime}(\phi_{T}),\psi_{T}\right\rangle_{\Gamma}\right)$
@@
$\quad-\varepsilon^{-1}\theta_{0}\left(\left\langle\phi^{n},\psi\right\rangle_{\Omega}+\frac{2}{h}\left\langle\phi_{B}^{n},\psi_{B}\right\rangle_{\Gamma}+\frac{2}{h}\left\langle\phi_{T}^{n},\psi_{T}\right\rangle_{\Gamma}\right).$ (3.17)
@@
$\quad+\frac{2}{hs}(-\Delta^{x}_{h})^{-1}(\phi_{B}-\phi^{n}_{B})_{i}+\frac{1}{h}\varepsilon^{-1}\left(I^{\prime}(\phi_{B,i})-\theta_{0}\phi^{n}_{B,i}\right)-\frac{1}{h}\kappa\Delta_{h}^{x}\phi_{B,i}+\tilde{C}_{B,n},$ (3.20)
@@
$\left.\frac{\mathrm{d}^{2}F_{h}^{n}}{\mathrm{d}\tau^{2}}(\phi+\tau\psi)\right|_{\tau=0}$ $=\frac{1}{s}\left\langle\psi,L_{h}^{-1}\psi\right\rangle_{\Omega}+\frac{2}{hs}\left\langle\psi_{B},(-\Delta_{h}^{x})^{-1}\psi_{B}\right\rangle_{\Gamma}+\frac{2}{hs}\left\langle\psi_{T},(-\Delta^{x}_{h})^{-1}\psi_{T}\right\rangle_{\Gamma}$
$\quad+\varepsilon^{-1}\left(\left\langle\psi I^{\prime\prime}(\phi),\psi\right\rangle_{\Omega}+\frac{2}{h}\left\langle\psi_{B}I^{\prime\prime}(\phi_{B}),\psi_{B}\right\rangle_{\Gamma}+\frac{2}{h}\left\langle\psi_{T}I^{\prime\prime}(\phi_{T}),\psi_{T}\right\rangle_{\Gamma}\right)$
@@
$=\frac{1}{s}\left\langle\psi,L_{h}^{-1}(\phi-\phi^{n})\right\rangle_{\Omega}+\frac{2}{hs}\left\langle\psi_{B},(-\Delta_{h}^{x})^{-1}(\phi_{B}-\phi^{n}_{B})\right\rangle_{\Gamma}+\frac{2}{hs}\left\langle\psi_{T},(-\Delta^{x}_{h})^{-1}(\phi_{T}-\phi^{n}_{T})\right\rangle_{\Gamma}$
$\quad+\varepsilon^{-1}\left(\left\langle I^{\prime}(\phi),\psi\right\rangle_{\Omega}+\frac{2}{h}\left\langle I^{\prime}(\phi_{B}),\psi_{B}\right\rangle_{\Gamma}+\frac{2}{h}\left\langle I^{\prime}(\phi_{T}),\psi_{T}\right\rangle_{\Gamma}\right)$
@@
$\quad-\varepsilon^{-1}\theta_{0}\left(\left\langle\phi^{n},\psi\right\rangle_{\Omega}+\frac{2}{h}\left\langle\phi_{B}^{n},\psi_{B}\right\rangle_{\Gamma}+\frac{2}{h}\left\langle\phi_{T}^{n},\psi_{T}\right\rangle_{\Gamma}\right).$ (3.25)
@@
$\quad+\frac{2}{hs}\Big{(}(-\Delta_{h}^{x})^{-1}(\phi^{\star}_{B}-\phi^{n}_{B})_{i_{0}}-(-\Delta_{h}^{x})^{-1}(\phi^{\star}_{B}-\phi^{n}_{B})_{i_{1}}\Big{)}$
$\quad+\varepsilon^{-1}(\frac{1}{2}+2h^{-1})\Big{(}\ln(1+\phi^{\star}_{i_{0},0})-\ln(1-\phi^{\star}_{i_{0},0})\Big{)}$
$\quad+\varepsilon^{-1}(\frac{1}{2}+2h^{-1})\Big{(}-\ln(1+\phi^{\star}_{i_{1},0})+\ln(1-\phi^{\star}_{i_{1},0})\Big{)}$
@@
$\quad-\varepsilon^{-1}\theta_{0}(\frac{1}{2}+2h^{-1})(\phi^{n}_{i_{0},0}-\phi^{n}_{i_{1},0}).$ (3.36)
In comparison with (3.27), in which the minimum value point $(i_{0},j_{0})$ is located within the interior of the grid, a distinguishing feature of the directional derivative (3.36) is the coefficient in front of the singular logarithmic terms, which is $\varepsilon^{-1}(\frac{1}{2}+2h^{-1})$, instead of $\varepsilon^{-1}$ in (3.27). In either case, the corresponding coefficients are positive, and this crucial fact will play an essential role in the positivity-preserving analysis.
@@
$\frac{1/2+2/h}{\varepsilon}\left(\ln(1+\phi^{\star}_{i_{0},0})-\ln(1+\phi^{\star}_{i_{0},0})-\ln(1-\phi^{\star}_{i_{0},0})+\ln(1-\phi^{\star}_{i_{1},0})\right)$ (3.37)
$\leq\frac{1/2+2/h}{\varepsilon}\left(\ln\left(\frac{\delta}{2-\delta}\right)+\ln\left(\frac{1-\beta_{B,0}}{1+\beta_{B,0}}\right)\right)$ (3.38)
@@
$-\frac{\theta_{0}(1+4/h)}{\varepsilon}\leq\frac{\theta_{0}(1/2+2/h)}{\varepsilon}\left(\phi^{n}_{i_{0},0}-\phi^{n}_{i_{1},0}\right)\leq\frac{\theta_{0}(1+4/h)}{\varepsilon}.$ (3.43)
@@
$\frac{1}{h^{2}}\mathrm{d}_{\tau}\!\left.F_{h}^{n}(\phi^{\star}+\tau\psi)\right|_{\tau=0}\leq\frac{1/2+2/h}{\varepsilon}\ln\left(\frac{\delta}{2+\delta}\right)+C_{4},$ (3.44)
@@
$C_{4}:=\frac{1/2+2/h}{\varepsilon}\ln\left(\frac{1-\beta_{B,0}}{1+\beta_{B,0}}\right)+\frac{C_{1}+4C_{2}/h}{s}+\frac{2\varepsilon}{h^{2}}+\theta_{0}\frac{1+4/h}{\varepsilon}.$
@@
$\frac{1}{h^{2}}\mathrm{d}_{\tau}\!\left.F_{h}^{n}(\phi^{\star}+\tau\psi)\right|_{\tau=0}\leq\frac{1/2+2/h}{\varepsilon}\ln\left(\frac{\delta}{2+\delta}\right)+C_{4}<0.$ (3.45)
@@
$\quad-\frac{2}{hs}(-\Delta_{h}^{x})^{-1}(\phi_{B}^{\star}-\phi_{B}^{n})_{i_{1}}+\varepsilon^{-1}\Big{(}\ln(1+\phi^{\star}_{i_{0},j_{0}})-\ln(1-\phi^{\star}_{i_{0},j_{0}})\Big{)}$
$\quad+\varepsilon^{-1}(\frac{1}{2}+2h^{-1})\Big{(}-\ln(1+\phi^{\star}_{i_{1},0})+\ln(1-\phi^{\star}_{i_{1},0})\Big{)}$
@@
$\quad-\varepsilon^{-1}\theta_{0}\phi^{n}_{i_{0},j_{0}}+\varepsilon^{-1}\theta_{0}(\frac{1}{2}+2h^{-1})\phi^{n}_{i_{1},0}.$ (3.46)
@@ 4 Total energy stability analysis
Remark 4.1.
Many studies have discussed the convergence of convex splitting schemes under classical periodical and homogeneous Neumann boundary conditions. Since the spatial truncation error, the exact solution is inconsistent with the discrete mass conservation. A Fourier projection (including sine or cosine projection) is commonly applied to treat the issue with acceptable artificial error [11, 12, 23]. For the semi-discrete convergence proposed in [5, 31], this problem does not happen. However, the Fourier projection is invalid for the dynamical boundary condition. A suitable alternative to the Fourier projection is needed, which is the difficulty in analyzing the convergence of the proposed scheme. We will consider this issue in our future research.
@@ 5 Numerical experiments
In this section, some numerical examples will be presented to validate the proposed scheme. A combination of multigrid method and Newton iteration is applied to deal with the nonlinear equations coupled between the interior region and the boundary section. A nonlinear full approximation scheme (FAS) multigrid method is applied for solving the semi-implicit numerical scheme (2.15) – (2.19). We present the convergence test and perform some 2-D sample computation results, while a 3-D extension could be similar. The computational domain is taken as $\Omega=(0,1)^{2}$, parameters will be specified when mentioned.